acid (HCI) reacts with solid sodium hydroxide (NaOH) to produce aqueous sodium chloride (NaCI) and liquid water (H,o), I3.31 g Aqueous of water is produced from the reaction of 12.4 g of hydrochloric acd and 24.5 g of sodium hydroxide, calculate the percent yield of water
Answer #1

HCl + NaOH --------> NaCl + H2O

36.5 g HCl reacts with 40 g NaOH to form 18 g water
12.4 g HCl requires 13.6 g NaOH ( We have 24.5 g NaOH) . HCl is the miting reactant and amount of H2O formed depends on the amount HCl

36.5 g HCl produces 18 g water
12.4 g HCl produces 12.4 g * 18 g / 36.5 g = 6.115 g H2O ( this is theoritical yield)
But actual yield of H2O is 3.31 g
% Yield of H2O = (3.31 / 6.115 )* 100 = 54.12 %

** If you like the explanation, it deserves your upvote/thumbsup :)